OpenStudy (anonymous):

how do you solve y=e^(x+2)

OpenStudy (mathmagician):

\[e ^{x}=y/(e ^{2}\]. \[x=\ln y+\ln(e ^{-2})=\ln y-2\]

OpenStudy (anonymous):

lny = lne^(x+2) lny = (x+2)lne lny = x+2 x = lny -2

OpenStudy (amistre64):

it has many solutions; you slove it by graphing the plots on a peice of paper ...